The molar volume of kal and nacl are 37.46 ml & 27.94 ml respectively. The ratio of the unit cube edges of the crystals is?

Grade:12

1 Answers

Vikas TU
7 years ago
From,
Vm = 4*pi*r1^3*Na/3
we get,
r1 = (3Vm/(4*pi*Na))^(1/3)
   = (3*37.46/4pi*Na)^(1/3)
and similarly for Nacl we get,
r2 = (3*27.94/4pi*Na)(1/3)
dividing both we get the ratio as:
r1/r2 = 1.105/1 or => 11:10